OpenStudy (pinterestaddict15):

18. The isotope, Carbon-14 emits beta radiation and has a half live of 5730 years. How long will it take for the amount of radiation this gives off to drop by a factor of 4?

OpenStudy (jiteshmeghwal9):

Amount of radioactive substance left after n half life periods is \[N=\left( \frac{1}{2} \right)^nN_0\] In ur question \(\Large{N=\frac{N_0}{4}}\)

OpenStudy (jiteshmeghwal9):

Also, \(n*t_{1/2}=T\) where T=total time taken in the radiation process

OpenStudy (jiteshmeghwal9):

\[N=\left( \frac{1}{2} \right)^n N_0\]\[\frac{N_0}{4}=\left( \frac{1}{2} \right)^nN_0\]so n=2 Total time = \(n*t_{1/2}\)

OpenStudy (osprey):

Is it 22, 920 years ? - two half lives 5730yrs per half life ?